OpenStudy (anonymous):

When the outliers are removed, how does the mean change? dot plot with 1 on 18, 2 on 38, 1 on 40, 3 on 44, 1 on 46, 1 on 48 The mean decreases by 2.75. The mean increases by 2.75. The mean increases by 3.25. There are no outliers.

OpenStudy (anonymous):

i know the mean increases

OpenStudy (anonymous):

by 3.25

OpenStudy (anonymous):

is it?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (misssunshinexxoxo):

I'll check this

OpenStudy (anonymous):

thank you

OpenStudy (misssunshinexxoxo):

Pretty sure by 3.25; I'll have a friend check @jim_thompson5910

OpenStudy (anonymous):

thats what i thought

OpenStudy (phi):

First , figure out 1 on 18, 2 on 38, 1 on 40, 3 on 44, 1 on 46, 1 on 48 18+2*38+40+3*44+46+48 and then divide by 1+2+1+3+1+1 that is the original mean then figure out 2*38+40+3*44+46+48 and divide by 2+1+3+1+1 that is the new mean

OpenStudy (anonymous):

its 42 and 1/3

OpenStudy (anonymous):

witch simplifys to 3.25

OpenStudy (phi):

What did you get for 18+2*38+40+3*44+46+48 ?

OpenStudy (anonymous):

360

OpenStudy (phi):

and what do you get for 1+2+1+3+1+1 ?

OpenStudy (anonymous):

9

OpenStudy (phi):

the average is the sum of all the numbers (i.e. 360) divided by the number of numbers (i.e. 9) so the first mean is 360/9 = what ?

OpenStudy (anonymous):

40

OpenStudy (phi):

now we "toss out the outlier" which is the number that seems "far away" from the others. in this case, the 18 seems far away so now we add up all the numbers 360 take away 18 we get 360-18 = 342 and instead of 9 numbers, we now have 8 numbers. the new mean is 342/8 = what ?

OpenStudy (anonymous):

42.75

OpenStudy (phi):

so old mean was 40, new mean is 42.75 the new mean is bigger by how much ?

OpenStudy (anonymous):

2.75

OpenStudy (phi):

and that is choice B